OpenStudy (anonymous):

Will Fan and Medal! Thank you♥ Write a rule for the function represented by the table. Function Table Input (x) - 0, 1, 2, 3, 4 Output (y) - 3, 6, 9, 12, 15 a. y= 3x b. y= 3x + 3 c. y= 5x + 1 d. y= x + 5

OpenStudy (mathmale):

"Write a rule for the function represented by the table. Function Table Input (x) - 0, 1, 2, 3, 4 Output (y) - 3, 6, 9, 12, 15 " This appears to be a linear function, with form y = mx + b. According to the input-output table, if x=1, y = 6 => (1,6) If x=4, y =15. => (4,15) How would you find the equation of a line that passes thru these 2 points? Hint: You need to find the slope first. Then, use y=mx + b to find b, or use the point-slope formula,\[y-y _{1}=m(x-x _{1})\]

OpenStudy (mathmale):

It's reasonable to expect that you know these equations for a straight line and how to use them.